40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 14 02:33:37 2016, ericq, Update, LSC, ALSX Noise still anomalously high  2016-01-14_ALSXspectra.pdfPDHsig.pdf
    Reply  Tue Jan 19 17:54:39 2016, gautam, Update, LSC, ALSX Noise still anomalously high  2016_01_19_ALS_OutOfLoop.pdfGreen_Locking_Trends.png
Message ID: 11937     Entry time: Tue Jan 19 17:54:39 2016     In reply to: 11931
Author: gautam 
Type: Update 
Category: LSC 
Subject: ALSX Noise still anomalously high  

While carrying out my end-table power investigations, I decided to take a quick look at the out-of-loop ALSX noise - see the attached plot. The feature at ~1kHz seems less prominent (factor of 2?) now, though its still present, and the overall noise above a few tens of Hz is still much higher than the reference. The green transmission was maximized to ~0.19 before this spectrum was taken.

EDIT 1130pm: 

We managed to access the trends for the green reflected and transmitted powers from a couple of months back when things were in their nominal state - see Attachment #2 for the situation then. For the X arm, the green reflected power has gone down from ~1300 counts (November 2015) to ~600 counts (january 2016) when locked to the arm and alignment is optimized. The corresponding numbers for the green transmitted powers (PSL + End Laser) are 0.47 (November 2015) and ~0.18 (January 2016). This seems to be a pretty dramatic change over just two months. For the Y-arm, the numbers are: ~3500 counts (Green REFL, Nov 2015), ~3500 counts (Green REFL, Jan 2016) ~1.3 (Green Trans, Nov 2015), ~1 (Green Trans, Jan 2016). So it definitely looks like something has changed dramatically with the X-end setup, while the Y-end seems consistent with what we had a couple of months ago...

Attachment 1: 2016_01_19_ALS_OutOfLoop.pdf  57 kB  | Show | Hide all | Show all
Attachment 2: Green_Locking_Trends.png  35 kB  Uploaded Wed Jan 20 00:18:48 2016  | Hide | Hide all | Show all
Green_Locking_Trends.png
ELOG V3.1.3-